# The cost of ensembling: is it always worth combining?

### v1

@@ Abstract
Given the continuous increase in dataset sizes and the complexity of forecasting models, the trade-off between forecast accuracy and computational cost is emerging as an extremely relevant topic, especially in the context of ensemble learning for time series forecasting. To asses it, we evaluated ten base models and eight ensemble configurations across two large-scale retail datasets (M5 and VN1), considering both point and probabilistic accuracy under varying retraining frequencies. We showed that ensembles consistently improve forecasting performance, particularly in probabilistic settings. However, these gains come at a substantial computational cost, especially for larger, accuracy-driven ensembles. We found that reducing retraining frequency significantly lowers costs, with minimal impact on accuracy, particularly for point forecasts. Moreover, efficiency-driven ensembles offer a strong balance, achieving competitive accuracy with considerably lower costs compared to accuracy-optimized combinations. Most importantly, small ensembles of two or three models are often sufficient to achieve near-optimal results. These findings provide practical guidelines for deploying scalable and cost-efficient forecasting systems, supporting the broader goals of sustainable AI in forecasting. Overall, this work shows that careful ensemble design and retraining strategy selection can yield accurate, robust, and cost-effective forecasts suitable for real-world applications.
@@ 1 Introduction
Ensemble learning has become a powerful and widely adopted approach in time series forecasting due to its ability to improve predictive accuracy and robustness by aggregating the outputs of multiple models. The fundamental idea behind ensemble methods is that combining forecasts from diverse models can help balance their individual strengths and weaknesses, thereby reducing the risk of overfitting and mitigating model-specific biases or errors. This approach leverages the principle that while individual models might make different types of errors, their combination can average out these errors. One of the most intriguing and empirically validated findings in the literature is that simple ensemble strategies, such as taking the mean or median of forecasts, often outperform more sophisticated combination methods like stacking (also known as meta-learning) or weighted ensembles. This phenomenon, known as the forecast combination puzzle (Claeskens et al., 2016), suggests that in many cases the added complexity of estimating optimal weights or training a meta-learner does not yield meaningful gains in accuracy and can even lead to overfitting, especially when the number of forecast points is limited. Indeed, the success of simple combinations is partly due to their ability to remain robust in the presence of model misspecification, parameter uncertainty, and changing data dynamics, issues that more complex ensemble methods may struggle to accommodate effectively.
In general, the benefits of ensemble learning are numerous: it improves forecast accuracy, enhances model stability across time and datasets, and increases resilience to structural breaks or shifts in the data. Ensembles also provide a practical hedge against model uncertainty, allowing practitioners to rely less on the assumption that any single model is correctly specified. In the context of global forecasting models, ensemble methods become especially valuable. Global models are trained across many time series simultaneously, and while they excel at identifying cross-series patterns, they can also suffer from model instability or biases if the chosen architecture fails to generalize across heterogeneous series. Combining multiple global models, each with different biases, architectures, or feature representations, can help counteract these issues by smoothing out individual model errors and reducing the risk of overfitting to shared but spurious signals across series (Wang et al., 2023). Additionally, ensembles serve as a hedge against model selection uncertainty, which is particularly relevant in global settings where the optimal model may vary substantially across groups of series. By pooling predictions, ensembles offer a form of model diversification that improves generalization, making them a powerful and practical tool in modern forecasting systems.
@@
Therefore, one of the major advantages of adopting a global modeling approach can be offset when ensemble learning is introduced. Combining several models, even with simple techniques like averaging, inherently increases the computational time, as multiple models must be trained, validated, and maintained, producing substantial computational overhead. In this sense, ensemble learning may weaken one of the core benefits of the global modeling paradigm: reduced training time and resource usage. Moreover, since forecasts are typically generated using cloud computing services with pay-as-you-go pricing models (Fotios Petropoulos & Spiliotis, 2024), increased computational time and resource consumption directly leads to higher forecasting costs for organizations. These costs may also rise exponentially when frequent retraining strategies are adopted. Indeed, it is a widespread practice to update forecasting models whenever new data becomes available, often driven by the belief that frequent updates enhance the model’s ability to adapt to evolving patterns and improve predictive accuracy. However, when ensembles are used, all the models within the ensembles must be retrained to obtain the final predictions, abruptly increasing the costs of forecasting.
This trade-off between performance and computational time in the context of ensemble learning raises important questions about the balance between accuracy gains and cost efficiency, especially in large-scale forecasting applications. However, despite these challenges, ensemble learning remains one of the most widely used strategies in the forecasting industry, which further amplifies concerns about the long-term sustainability and scalability of forecasting systems. Indeed, from a practical perspective, forecast combinations, coupled with frequent retraining, has also significant environmental costs, since the energy consumption of model training extends beyond direct computational costs, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Therefore, understanding the cost of ensemble forecasting models and the effects of retraining on their performance is of paramount importance for advancing more sustainable forecasting practices.
@@ 1.1 Research Question
We aim to address the question ”Are ensembles always the solution?”. Specifically, we focus on the trade-off between accuracy and sustainability, in terms of the cost of producing the forecasts, among different types of ensembles of global forecasting models. This cost is particularly relevant for ensemble models since it becomes exponentially larger with the number of base models used to produce the combined forecasts. To this end, we use ten global models as base learners, ranging from traditional machine learning methods to advanced deep learning architectures. In addition, we evaluate eight ensemble approaches, all trained and tested on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition datasets.
We also investigated the effects of retraining on the forecasting performance of ensembles, compared to that of the base models. In particular, we assess whether reducing the frequency of retraining, by avoiding re-estimation of base models as new data becomes available, can effectively enhance the cost-accuracy trade-off in ensemble forecasting. Hence, we explored a wide range of retraining strategies, from continuous retraining to a no-retraining approach, including various intermediate periodic retraining schemes to encompass the most practical and effective scenarios.
@@ 1.2 Contributions
    We analyze different ensemble strategies to asses the effects of the ensemble size on forecast accuracy and costs.
@@ 2 Related works
The cross-learning approach in time series forecasting has witnessed substantial advancements in recent years, becoming a central theme in contemporary research. Today, global models are commonly used as benchmarks in empirical studies, underscoring their growing relevance in the field Semenoglou et al. (2021). Additionally, theoretical contributions from Montero-Manso & Hyndman (2021) and Montero-Manso (2023) have shown that global models can achieve accuracy comparable to local models with reduced complexity and without relying on assumptions about data similarity. Indeed, global models have proven particularly effective across a variety of forecasting domains, especially in retail forecasting domain (Makridakis et al., 2022a), and several methods have been proposed to further boost their performance (Godahewa et al., 2021a; Bandara et al., 2020, 2021). In particular, ensemble learning is a powerful tool in the forecasters’ hands offering a reliable strategy to enhance accuracy and robustness by combining the strengths of multiple models, not only in the cross-learning context (Wang et al., 2023). Relative to point forecasting, many different combination methods have been proposed: linear combinations with optimal weights (Newbold & Granger, 1974), performance based weighting (Pawlikowski & Chorowska, 2020), criteria-based weighting (Kolassa, 2011), and many different approaches based on stacking, or meta-learning (Wang et al., 2023). However, time series forecasting competitions have shown that simple ensemble strategies, like simple mean or median, are extremely accurate, yet efficient (Makridakis et al., 2022b). Indeed, simple combination schemes are hard to beat, and the simple arithmetic average of predictions with equal weights remains the most widely used and surprisingly effective combination rule (Claeskens et al., 2016).
@@
Most evaluations of global models emphasize point forecast accuracy, likely due to the fact that many machine learning and deep learning algorithms do not natively produce probabilistic outputs (Makridakis et al., 2022c). However, in applications such as supply chain management, quantifying uncertainty is essential, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). In this regard, the Conformal Inference framework introduced by Vovk et al. (2005) offers a flexible tool for uncertainty quantification, and has been successfully applied to time series forecasting problems (Stankeviciute et al., 2021). Forecast combinations in the context of probabilistic predictions are an active area of research. Different methods have been proposed, from linear pooling, Bayesian model averages, and integral combinations (Wang et al., 2023). However, quantile aggregation via simple average has proven to be effective and accurate (Smyl & Hua, 2019; Busetti, 2017).
@@
From the perspective of model retraining, the most comprehensive investigation in the context of global models has been conducted by Zanotti (2025), who found that reducing the retraining frequency of global models can lower forecasting costs without harming accuracy. Other works, such as Spiliotis & Petropoulos (2024), explored retraining strategies and parameter updates for local exponential smoothing models, while Huber & Stuckenschmidt (2020) examined retraining in retail demand forecasting using a limited range of models and datasets. Although these studies offer useful insights, direct analysis on the effects of retraining ensemble models remains unexplored.
@@
The question of whether ensemble methods are consistently worthwhile is relevant both for the forecasting and the broader machine learning community. Indeed, the growing advocacy for sustainable AI (Schwartz et al., 2020; Getzner et al., 2023) highlights the importance of evaluating the environmental and computational costs of model development and maintenance. Our study directly addresses this gap by systematically evaluating different ensemble strategies’ cost-accuracy trade-off. We also assess how different updating frequencies affect the forecasting accuracy across the ensemble models, with the aim of testing whether less frequent retraining can be an effective tool to manage this trade-off. Answering these questions, this research provides both theoretical clarity and practical guidelines towards more sustainable forecasting practices.
@@ 3.1 Datasets
For our experiments, we utilized two prominent retail forecasting datasets: the M5 and VN1 competition datasets. The M5 competition, part of the M-competitions series led by Spyros Makridakis and collaborators, was designed to benchmark forecasting methods in a retail demand setting (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) is widely recognized and extensively studied, comprising 3,049 daily time series representing unit sales of Walmart products. These sales span three main product categories, Food, Hobbies, and Household, across ten stores in three U.S. states: California, Texas, and Wisconsin. Covering the period from 2011 to 2016, the dataset features highly intermittent time series organized hierarchically, enabling multi-level forecasting (e.g., individual SKUs, categories, stores, and states). It also includes relevant exogenous variables such as prices, promotions, and special events like holidays. The VN1 Forecasting – Accuracy Challenge, launched in October 2024 by Flieber, Syrup Tech, and SupChains, marked the first edition of its kind (Vandeput, 2024). This dataset consists of weekly sales data for 15,053 products sold from 2020 to 2024 by various e-vendors, primarily based in the U.S. Unlike the M5 dataset, which contains sales from a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset aggregates sales across 328 warehouses operated by 46 distinct retailers. To the best of our knowledge, we are among the first to benchmark forecasting models on this dataset. Together, these two datasets represent the most recent and comprehensive publicly available collections of time series related to retail demand, providing a strong foundation for generalizing our findings in the domain of demand forecasting.
Table 1: The M5 and the VN1 datasets used in the experiments.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
In both datasets we concentrate on the most granular level (individual SKUs), because the potential gains from reduced retraining are greatest at lower aggregation levels. To ensure a consistent evaluation strategy (Section 3.3), we filtered the series: daily SKUs retained at least two years of data (more than 730 observations), while weekly SKUs required a minimum of three years (more than 157 observations).

### v2

@@ Abstract
Given the continuous increase in dataset sizes and the complexity of forecasting models, the trade-off between forecast accuracy and computational cost is emerging as an extremely relevant topic, especially in the context of ensemble learning for time series forecasting. To assess it, we evaluated ten base models and eight ensemble configurations across two large-scale retail datasets, considering both point and probabilistic accuracy under varying retraining frequencies. We showed that ensembles consistently improve forecasting performance, particularly in probabilistic settings. However, these gains come at a substantial computational cost, especially for larger, accuracy-driven ensembles. We found that reducing retraining frequency significantly lowers costs, with minimal impact on accuracy, particularly for point forecasts. Moreover, efficiency-driven ensembles offer a strong balance, achieving competitive accuracy with considerably lower costs compared to accuracy-optimized combinations. Most importantly, small ensembles of two or three models are often sufficient to achieve near-optimal results. These findings provide practical guidelines for deploying scalable and cost-efficient forecasting systems, supporting the broader goals of sustainable AI in forecasting. Overall, this work shows that careful ensemble design and retraining strategy selection can yield accurate, robust, and cost-effective forecasts suitable for real-world applications.
@@ 1 Introduction
Ensemble learning has become a powerful and widely adopted approach in time series forecasting due to its ability to improve predictive accuracy and robustness by aggregating the outputs of multiple models. The fundamental idea behind ensemble methods is that combining forecasts from diverse models can help balance their individual strengths and weaknesses, thereby reducing the risk of overfitting and mitigating model-specific biases or errors. This approach leverages the principle that while individual models might make different types of errors, their combination can average out these errors. One of the most intriguing and empirically validated findings in the literature is that simple ensemble strategies, such as taking the mean or median of forecasts, often outperform more sophisticated combination methods like stacking (also known as meta-learning) or weighted ensembles. This phenomenon, known as the forecast combination puzzle (Claeskens et al., 2016), suggests that in many cases, the added complexity of estimating optimal weights or training a meta-learner does not yield meaningful gains in accuracy and can even lead to overfitting, especially when the number of forecast points is limited. Indeed, the success of simple combinations is partly due to their ability to remain robust in the presence of model misspecification, parameter uncertainty, and changing data dynamics, issues that more complex ensemble methods may struggle to accommodate effectively.
In general, the benefits of ensemble learning are numerous: it improves the forecast accuracy, enhances the model stability across time and datasets, and increases resilience to structural breaks or shifts in the data. Ensembles also provide a practical hedge against model uncertainty, allowing practitioners to rely less on the assumption that any single model is correctly specified. In the context of global forecasting models, ensemble methods become especially valuable. Global models are trained across many time series simultaneously, and while they excel at identifying cross-series patterns, they can also suffer from model instability or biases if the chosen architecture fails to generalize across heterogeneous series. Combining multiple global models, each with different biases, architectures, or feature representations, can help counteract these issues by smoothing out individual model errors and reducing the risk of overfitting to shared but spurious signals across series (Wang et al., 2023). Additionally, ensembles serve as a hedge against model selection uncertainty, which is particularly relevant in global settings where the optimal model may vary substantially across groups of series. By pooling predictions, ensembles offer a form of model diversification that improves generalization, making them a powerful and practical tool in modern forecasting systems.
@@
Therefore, one of the major advantages of adopting a global modeling approach, namely scalability, can be offset when ensemble learning is introduced in the forecasting system. Combining several models, even with simple techniques like averaging, inherently increases the computational time, as multiple models must be trained, validated, and maintained, producing substantial computational overhead. In this sense, ensemble learning may weaken one of the core benefits of the global modeling paradigm: reduced training time and resource usage. Moreover, since forecasts are typically generated using cloud computing services with pay-as-you-go pricing models (Fotios Petropoulos & Spiliotis, 2024), increased computational time and resource consumption directly leads to higher forecasting costs for organizations. These costs may also rise exponentially when frequent retraining strategies are adopted. Indeed, it is a widespread practice to update forecasting models whenever new data becomes available, often driven by the belief that frequent updates enhance the model’s ability to adapt to evolving patterns and improve predictive accuracy. However, when ensembles are used, all the models within the ensembles must be retrained to obtain the final predictions, abruptly increasing the costs of forecasting.
This trade-off between performance and computational time in the context of ensemble learning raises important questions about the balance between accuracy gains and cost efficiency, especially in large-scale forecasting applications. However, despite these challenges, ensemble learning remains one of the most widely used strategies in the forecasting industry, which further amplifies concerns about the long-term sustainability and scalability of forecasting systems. Indeed, from a practical perspective, forecast combinations, coupled with frequent retraining, has also significant environmental costs, since the energy consumption of model training extends beyond direct computational costs, contributing to global energy consumption and carbon emissions (Schwartz et al., 2020). Therefore, understanding the cost of ensembling forecasting models and the effects of retraining on their performance are of paramount importance for advancing more sustainable forecasting practices.
@@ 1.1 Research Question
We aim to address the question ”Are ensembles always the solution?”. Specifically, we focus on the trade-off between accuracy and sustainability, in terms of the cost of producing the forecasts, among different types of ensembles of global forecasting models. This cost is particularly relevant for ensemble models since it becomes exponentially larger with the number of base models used to produce the combined forecasts. To this end, we use ten global models as base learners, ranging from traditional machine learning methods to advanced deep learning architectures. In addition, we simulate eight ensemble approaches, all trained and tested on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition datasets.
We also investigate the effects of retraining on the forecasting performance of ensembles, compared to that of the base models. In particular, we assess whether reducing the frequency of retraining, by avoiding re-estimation of base models as new data becomes available, can effectively enhance the cost-accuracy trade-off in ensemble forecasting. Hence, we explore a wide range of retraining strategies, from continuous retraining to a no-retraining approach, including various intermediate periodic retraining schemes to encompass the most practical and effective scenarios.
@@ 1.2 Contributions
    We analyze different ensemble strategies to assess the effects of the ensemble size on forecast accuracy and costs.
@@ 2 Related works
The cross-learning (global) approach in time series forecasting has witnessed substantial advancements in recent years becoming a central theme in contemporary research. Today, global models are commonly used as benchmarks in empirical studies, underscoring their growing relevance in the field (Semenoglou et al., 2021). Additionally, theoretical contributions from Montero-Manso & Hyndman (2021) and Montero-Manso (2023) have shown that global models can achieve accuracy comparable to local models with reduced complexity and without relying on assumptions about data similarity. Indeed, global models have proven particularly effective across a variety of forecasting domains, especially in retail forecasting (Makridakis et al., 2022a), and several methods have been proposed to further boost their performance (Godahewa et al., 2021a; Bandara et al., 2020, 2021). In particular, ensemble learning is a powerful tool in the forecasters’ hands offering a reliable strategy to enhance accuracy and robustness by combining the strengths of multiple models, not only in the cross-learning context (Wang et al., 2023). Relative to point forecasting, many different combination methods have been proposed: linear combinations with optimal weights (Newbold & Granger, 1974), performance based weighting (Pawlikowski & Chorowska, 2020), criteria-based weighting (Kolassa, 2011), and many different approaches based on stacking, or meta-learning (Wang et al., 2023). However, time series forecasting competitions have shown that simple ensemble strategies, like simple mean or median, are extremely accurate, yet efficient (Makridakis et al., 2022b). Indeed, simple combination schemes are hard to beat, and the simple arithmetic average of predictions with equal weights remains the most widely used and surprisingly effective combination rule (Claeskens et al., 2016).
@@
Most evaluations of global models emphasize point forecast accuracy, likely due to the fact that many machine learning and deep learning algorithms do not natively produce probabilistic outputs (Makridakis et al., 2022c). However, in applications such as supply chain management, quantifying uncertainty is essential, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). Forecast combinations in the context of probabilistic predictions are an active area of research. Different methods have been proposed, from linear pooling, Bayesian model averages, and integral combinations (Wang et al., 2023). However, quantile aggregation via simple average has proven to be effective and accurate (Smyl & Hua, 2019; Busetti, 2017).
@@
From the perspective of model retraining, the most comprehensive investigation in the context of global models has been conducted by Zanotti (2025), who found that reducing the retraining frequency of global models can lower forecasting costs without harming accuracy. Other works, such as Spiliotis & Petropoulos (2024), explored retraining strategies and parameter updates for local exponential smoothing models, while Huber & Stuckenschmidt (2020) examined retraining in retail demand forecasting using a limited range of models and datasets. Although these studies offer useful insights, direct analysis on the effects of retraining on ensemble models remains unexplored.
@@
The question of whether ensemble methods are consistently worthwhile is relevant both for the forecasting and the broader machine learning community. Indeed, the growing advocacy for sustainable AI (Schwartz et al., 2020; Getzner et al., 2023) highlights the importance of evaluating the environmental and computational costs of model development and maintenance. Our study directly addresses this gap by systematically evaluating different ensemble strategies’ cost-accuracy trade-off. We also assess how different retraining frequencies affect the forecasting accuracy across the ensemble models, with the aim of testing whether less frequent retraining can be an effective tool to manage this trade-off. Answering these questions, this research provides both theoretical clarity and practical guidelines towards more sustainable forecasting practices.
@@ 3.1 Datasets
For our experiments, we utilized two prominent retail forecasting datasets: the M5 and VN1 competition datasets. The M5 competition, part of the M-competitions series led by Spyros Makridakis and collaborators, was designed to benchmark forecasting methods in a retail demand setting (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) is widely recognized and extensively studied, comprising 30,490 daily time series representing unit sales of Walmart products. These sales span three main product categories, Food, Hobbies, and Household, across ten stores in three U.S. states: California, Texas, and Wisconsin. Covering the period from 2011 to 2016, the dataset features highly intermittent time series organized hierarchically, enabling multi-level forecasting (e.g., individual SKUs, categories, stores, and states). It also includes relevant exogenous variables such as prices, promotions, and special events like holidays. The VN1 Forecasting – Accuracy Challenge, launched in October 2024 by Flieber, Syrup Tech, and SupChains, marked the first edition of its kind (Vandeput, 2024). This dataset consists of weekly sales data for 15,053 products sold from 2020 to 2024 by various e-vendors, primarily based in the U.S. Unlike the M5 dataset, which contains sales from a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset aggregates sales across 328 warehouses operated by 46 distinct retailers. To the best of our knowledge, we are among the first to benchmark ensemble forecasting models on this dataset. Together, these two datasets represent the most recent and comprehensive publicly available collections of time series related to retail demand, providing a strong foundation for generalizing our findings in the domain of demand forecasting.
In both datasets, we concentrated on the most disaggregated level (SKU $\times$ Store), because the potential gains from reduced retraining are greatest at lower aggregation levels. To ensure a consistent evaluation strategy (see Section 3.3), we filtered the series: daily SKUs retained at least two years of data (more than 730 observations), while weekly SKUs required a minimum of three years (more than 157 observations).